# Studying propagating turbulent structures in the near wake of a sphere using Hilbert proper orthogonal decomposition

### v1

@@ 4.2 Pairing POD Modes using HPOD Modes
Figure 8: Phase-averaged (a) streamwise and (c) transverse components of the ninth and tenth POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the fourth HPOD mode, adjusted for the phase angle offset $\Delta\phi_{4,9}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
The streamwise component of the phase-averaged ninth and tenth POD modes, shown in figure 8a, is qualitatively similar to the phase average of the fourth HPOD mode, shown in figure 8b, for $x/D>2.5$. In this domain, the streamwise components of both the POD modes and HPOD mode exhibit V-shaped structures similar to those in the fourth and fifth POD modes and the second HPOD mode. For $x/D<2.5$, there is more detail in the structure of the POD modes, which is most noticeable in figures 8a(iii) and 8a(iv) compared to figures 8b(iii) and 8b(iv). As the V-shaped structures in the streamwise components are more prominent in the phase-averaged ninth and tenth POD modes than the fourth HPOD mode, it can be inferred that they do not propagate in the streamwise direction.
@@
The transverse components of phase average of the ninth and tenth POD modes and the phase average of the fourth HPOD modes, shown in figures 8c and 8d, respectively, are also qualitatively similar. As is the case for the streamwise components, the transverse component of the phase-averaged POD modes contains features not present in the HPOD mode in the very near wake, due to the filtering effect of the Hilbert transform. This can be seen clearly in figures 8c(ii) and 8c(iii) compared to figures 8d(ii) and 8d(iii).
@@
Figure 9: Joint probability density function of the phase angles of (a) the first HPOD mode and the first and second POD modes, (b) the second HPOD mode and the fourth and fifth POD modes, (c) the third HPOD mode and the sixth and eighth POD modes, and (d) the fourth HPOD mode and the ninth and tenth POD modes. Phase angles are relative to the phase angle of each mode in the first instantaneous velocity field.
The joint probability density functions of the HPOD modes and the corresponding POD mode pairs, shown in figure 9, are all the greatest around $\phi^{a}_{k}=\phi_{i,j}$. The phase angles of both have been adjusted by the phase angle of the first instantaneous velocity field to account for the phase shift between the POD modes and the HPOD modes and to ensure all phase angles are in the domain $\phi\in[0,2\pi]$.
@@ 4.3 Pairing POD Modes using the Hilbert Transform
While HPOD modes can be used to determine matched pairs of POD modes, this process requires performing both decompositions. As the Hilbert transform can be applied in the streamwise direction, it can be applied directly to the POD modes despite their lack of temporal information. The result is a shift of the fundamental wavenumbers of each POD mode by $\mp 90^{\circ}$, which is analogous to a propagation of the mode downstream. Thus, the streamwise Hilbert transform of a POD mode is analogous to a POD mode which represents a $\mp 90^{\circ}$ phase shift of the same streamwise periodic structure.
In order to determine matched pairs of POD modes using the Hilbert transform, the Hilbert transform of each POD mode $\psi_{k}(\boldsymbol{x})$ in the streamwise direction $\mathcal{H}_{x}[\psi_{k}(\boldsymbol{x})]$ is used in place of the imaginary part of the HPOD modes to define the analytic signal of $\psi_{k}(\boldsymbol{x})$
@@
$\mathcal{R}_{k^{a},j}=\frac{\operatorname{cov}\big{(}(\psi_{k})^{a}(\boldsymbol{x},\Delta\phi_{k,j}),\psi_{j}(\boldsymbol{x})\big{)}}{\sqrt{\operatorname{var}\big{(}(\psi_{k})^{a}(\boldsymbol{x},\Delta\phi_{k,j})\big{)}\operatorname{var}\big{(}\psi_{j}(\boldsymbol{x})\big{)}}}.$ (49)
@@
The $R^{2}$ value of the phase average of the $k^{th}$ and $j^{th}$ POD modes compared to the phase average of the analytic signal of the $k^{th}$ POD mode is given by
$R^{2}_{k^{a},\widetilde{(k,j)}}=1-\frac{\sum\big{(}(\psi_{k})^{a}(\boldsymbol{x},\phi)-\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)\big{)}^{2}}{\sum\big{(}\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)\big{)}^{2}},$ (50)
@@
where $(\psi_{k})^{a}(\boldsymbol{x},\phi)$ is the analytic signal of the $k^{th}$ POD mode and $\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)$ is the phase average of the $k^{th}$ and $j^{th}$ POD modes at phase angle $\phi$, which was determined using phase angles $\phi=[0,2\pi]$ in increments of $0.01\pi$. The POD mode which best matches the analytic signal of the each of the first ten POD modes, the corresponding correlation coefficient, and the $R^{2}$ of the phase averages are given in table 5.
POD Mode $k$ Paired POD Mode $j$ $\Delta\phi_{k,j}$ $\mathcal{R}_{k^{a},j}$ $R^{2}_{k^{a},\widetilde{(k,j)}}$
1 2 $-0.5\pi$ 0.88 0.89
2 1 $0.5\pi$ 0.76 0.76
3 2 $0.5\pi$ 0.70 0.74
4 5 $-0.5\pi$ 0.79 0.81
5 4 $0.5\pi$ 0.72 0.72
6 8 $-0.5\pi$ 0.80 0.82
7 9 $-0.5\pi$ 0.60 0.67
8 6 $0.5\pi$ 0.76 0.77
9 10 $-0.5\pi$ 0.64 0.65
10 9 $0.5\pi$ 0.63 0.65
@@
Table 5: Paired POD modes using the analytic signal of the $k^{th}$ POD mode, and the corresponding phase angle, correlation coefficient and $R^{2}$ of the phase averages.
@@
For each of the first ten POD modes, the paired POD mode corresponds to the analytic signal at a phase angle of $\mp 0.5\pi$. Thus, the $j^{th}$ POD mode best matches the Hilbert transform of the corresponding $k^{th}$ POD mode. A pair of POD modes representing a structure propagating in the streamwise direction therefore requires that the POD modes are a mutual best match with phase angle shifts of $0.5\pi$ with opposite signs. Thus, the POD mode pairs identified using the Hilbert transform are the first and second POD modes, the fourth and fifth POD modes, the sixth and eighth POD modes, and the ninth and tenth POD modes. These are the same pairings as those identified using the HPOD modes.
@@
Figure 10: Phase-averaged (a) streamwise and (c) transverse components of the first and second POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the first POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
The streamwise component of the analytic signal of the first POD mode, shown in figure 10b, is a close match to the phase average of the first two POD modes, shown in figure 10a. Both show the same long regions of alternating sign propagating downstream from the sphere, with the main variation between them being for $x/D<2$, where the non-propagating features are located, as shown in figure 5a(iv). This is an artefact of applying the Hilbert transform to the POD mode, due to these structures being present in the second POD mode but not the first POD mode, on which the Hilbert transform was performed. This also explains the variation in the $R^{2}$ of the analytic signal of the first POD mode and the second POD mode and that of the analytic signal of the second POD mode and the first POD mode, being 0.89 and 0.76, respectively. This difference is due to the non-propagating components of these modes being absent in the other mode, but propagated by the Hilbert transform.
The variation between the transverse components of the phase-averaged first and second POD modes, shown in figure 10c, and the analytic signal of the first POD mode, shown in figure 10d, is significantly less than in the streamwise components. Both exhibit the same symmetric regions of alternating sign, which increase in length and decrease in intensity as they travel downstream. This is consistent with the vertical fluctuations from the flow coming around the sphere decaying once they are sufficiently far into the wake.
@@
Figure 11: Phase-averaged (a) streamwise and (c) transverse components of the fourth and fifth POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the fourth POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
@@
The streamwise component of the analytic signal of the fourth POD mode, shown in figure 11b, retains more of the asymmetry of the phase-averaged fourth and fifth POD modes, shown in figure 11a, than the second HPOD mode, shown in figure 6b. This is also true for the transverse component of the analytic signal of the fourth POD mode, shown in figure 11d. The transverse component of the analytic signal of the fourth POD mode, shown in figure 11d, also retains the asymmetry seen in the phase average of the fourth and fifth POD modes, shown in figure 11c. Although the filtering of the Hilbert transform is less prominent in the analytic signal of the fourth POD mode than in the second HPOD mode, shown in figures 6b and 6d, it is still present, most noticeably figures 11b(iii) and 11d(iv).
@@
Figure 12: Phase-averaged (a) streamwise and (c) transverse components of the sixth and eighth POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the sixth POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
While there is some filtering in the analytic signal of the sixth POD mode, the asymmetry seen in the phase average of the sixth and eighth POD modes is preserved where the HPOD modes were planar symmetric. This can be seen in the positive region introduced to the streamwise component of the third HPOD mode at $x/D=2.5$, shown in figure 7b(i), which is not present in the phase average of the sixth and eight POD modes, shown in figure 12a(i), or the analytic signal of the sixth POD mode, shown in figure 12b(i). It is also shown in the asymmetric region at $x/D=3.5$ in the transverse component of the sixth and eighth POD modes, shown in figure 12c(i), which is also off-centre in the analytic signal of the sixth POD mode, shown in figure 12d(i), but located at the centreline in the third HPOD mode, shown in figure 7d(i).
@@
Additionally, the intensity of the structures further downstream, such as in figure 12a(i), is clearly preserved in the analytic signal of the sixth POD mode, as shown in figure 12b(i), where the intensity of the same region is significantly reduced in the third HPOD mode, shown in figure 7b(i). This intensity in the phase average of the sixth and eighth POD modes around $x/D=4$, shown in figure 12a(v), is propagated downstream in the analytic signal of the sixth mode, as shown in figure 12b(v), whereas the intensity is lower in the same region in the third HPOD mode, as shown in figure 7b(v). This shows that the filtering which occurs from the Hilbert transform also affects the analytic signal of the POD modes.
@@
Figure 13: Phase-averaged (a) streamwise and (c) transverse components of the ninth and tenth POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the ninth POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$.
@@
While there is a significant difference between the streamwise components of the analytic signal of the ninth POD mode, shown in figure 13b, and that of the phase average of the ninth and tenth POD modes, shown in figure 13a, for $x/D<2$, they are more similar further downstream of the sphere. The structures in the phase average of the ninth and tenth modes are smaller scale for $x/D<2$ than they are for $x/D>2$, which is consistent with smaller scale fluctuations in the near wake breaking down as they travel downstream while the larger scales remain. As a result, these scales are not propagated by the Hilbert transform and are thus filtered in the analytic signal. The differences between the transverse components of the analytic signal of the ninth POD mode, shown in figure 13d, and the phase average of the ninth and tenth POD modes, shown in figure 13c, are significantly less prominent. However, there is still variation in the structures upstream of $x/D=1.5$, as shown in figure 13c(ii) and 13d(ii), and figures 13c(iii) and 13d(iii).

### v2

@@ 4.2 Pairing POD Modes using HPOD Modes
Figure 9: Phase-averaged (a) streamwise and (c) transverse components of the ninth and tenth POD modes, and the phase-averaged (b) streamwise and (d) transverse components of the fourth HPOD mode, adjusted for the phase angle offset $\Delta\phi_{4,9}$, at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 4 for an animated version of this figure.
@@
The streamwise component of the phase-averaged ninth and tenth POD modes, shown in figure 9a, is qualitatively comparable to the phase average of the fourth HPOD mode, shown in figure 9b, for $x/D>2.5$. In this domain, the streamwise components of the POD and HPOD modes exhibit V-shaped structures similar to those in the fourth and fifth POD modes and the second HPOD mode. For $x/D<2.5$, the structure of the POD modes exhibits greater detail, most evident in figures 9a(iii) and 9a(iv) compared with figures 9b(iii) and 9b(iv). The transverse components of the phase average of the ninth and tenth POD modes and the phase average of the fourth HPOD mode, shown in figures 9c and 9d, respectively, also exhibit qualitative similarity. Similar to the streamwise components, the transverse component of the phase-averaged POD modes includes features absent from the HPOD mode in the immediate near wake, attributable to the assumed periodicity of the Hilbert transform. This is clearly evident in figures 9c(ii) and 9c(iii) compared to figures 9d(ii) and 9d(iii) around $x/D=1$.
@@
Figure 10: Joint probability density function of the phase angles of (a) the first HPOD mode and the first and second POD modes, (b) the second HPOD mode and the fourth and fifth POD modes, (c) the third HPOD mode and the sixth and eighth POD modes, and (d) the fourth HPOD mode and the ninth and tenth POD modes. Phase angles of the HPOD modes are relative to the phase offset to the first POD mode of the corresponding pair.
The joint probability density functions of the HPOD modes and the corresponding POD mode pairs, shown in figure 10, are highest around $\phi^{a}_{k}=\phi_{i,j}$ when adjusted for the appropriate phase offset. This indicates that the phase relationship between each HPOD mode and its corresponding POD mode pair is consistent throughout the measurement domain. The spread of the distributions varies between the different mode pairs, with the tightest clustering occurring for the first HPOD mode and the first and second POD modes, shown in figure 10a, and the widest spread being that of the third HPOD mode and the sixth and eighth POD modes, shown in figure 10c. This is consistent with the $R^{2}_{k,\widetilde{(i,j)}}$ values of the mode shapes summarised in table 4, with the highest value of 0.83 occurring for the first HPOD mode and the first and second POD modes, and the lowest value of 0.67 occurring for the third HPOD mode and the sixth and eighth POD modes.
@@ 4.3 Pairing POD Modes using the Hilbert Transform
While HPOD modes can be used to identify matched pairs of POD modes, this process requires performing both decompositions. Because the Hilbert transform can be applied in the streamwise direction, it can be applied directly to the POD modes, despite their lack of temporal information. The result is a $\mp\pi/2$ phase shift of the fundamental Fourier components of each POD mode, analogous to the downstream propagation of the mode. Thus, the streamwise Hilbert transform of a POD mode is analogous to a POD mode that represents a $\mp\pi/2$ phase shift of the same streamwise periodic structure.
In order to determine matched pairs of POD modes using the Hilbert transform, the Hilbert transform of the $k^{th}$ POD mode in the streamwise direction $\mathcal{H}_{x}[\psi_{k}(\boldsymbol{x})]$ is used to define the analytic signal of $\psi_{k}(\boldsymbol{x})$
@@
$\mathcal{R}_{k^{a},j}(\Delta\phi_{k,j})=\frac{\operatorname{cov}\big((\psi_{k})^{a}(\boldsymbol{x},\Delta\phi_{k,j}),\psi_{j}(\boldsymbol{x})\big)}{\sqrt{\operatorname{var}\big((\psi_{k})^{a}(\boldsymbol{x},\Delta\phi_{k,j})\big)\operatorname{var}\big(\psi_{j}(\boldsymbol{x})\big)}}.$ (51)
@@
The $R^{2}_{k^{a},\widetilde{(k,j)}}$ value of the phase average of the $k^{th}$ and $j^{th}$ POD modes compared to the phase average of the analytic signal of the $k^{th}$ POD mode is given by
$R^{2}_{k^{a},\widetilde{(k,j)}}=1-\frac{\sum_{\boldsymbol{x},\phi}\big((\psi_{k})^{a}(\boldsymbol{x},\phi)-\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)\big)^{2}}{\sum_{\boldsymbol{x},\phi}\big(\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)\big)^{2}},$ (52)
@@
where $\boldsymbol{x}$ denotes the spatial coordinates, $(\psi_{k})^{a}(\boldsymbol{x},\phi)$ is the analytic signal of the $k^{th}$ POD mode and $\widetilde{\psi}_{k,j}(\boldsymbol{x},\phi)$ is the phase average of the $k^{th}$ and $j^{th}$ POD modes at phase angle $\phi$, which was determined using phase angles $\phi=[0,2\pi]$ in increments of $0.01\pi$. The POD mode which best matches the analytic signal of each of the first ten POD modes, is given in table 5, along with the corresponding correlation coefficient, and the $R^{2}_{k^{a},\widetilde{(k,j)}}$ of the phase averages. Both the correlation coefficients and $R^{2}_{k^{a},\widetilde{(k,j)}}$ values were calculated for the combined streamwise and transverse components of the modes, as well as the individual components.
POD Mode $k$ POD Mode $j$ $\Delta\phi_{k,j}$ $\mathcal{R}_{k^{a},j}$ $\mathcal{R}_{k^{a},j}^{u}$ $\mathcal{R}_{k^{a},j}^{v}$ $R^{2}_{k^{a},\widetilde{(k,j)}}$ $(R^{2}_{k^{a},\widetilde{(k,j)}})^{u}$ $(R^{2}_{k^{a},\widetilde{(k,j)}})^{v}$
1 2 $0.47\pi$ 0.85 0.82 0.89 0.86 0.82 0.88
2 1 $0.46\pi$ 0.78 0.75 0.82 0.78 0.74 0.81
3 2 $0.47\pi$ 0.77 0.75 0.79 0.79 0.77 0.80
4 5 $0.51\pi$ 0.86 0.84 0.89 0.87 0.84 0.89
5 4 $0.51\pi$ 0.79 0.76 0.82 0.79 0.75 0.82
6 8 $0.49\pi$ 0.81 0.73 0.86 0.83 0.77 0.87
7 9 $0.47\pi$ 0.59 0.61 0.58 0.67 0.67 0.67
8 6 $0.48\pi$ 0.73 0.63 0.82 0.73 0.61 0.80
9 10 $0.46\pi$ 0.75 0.71 0.82 0.75 0.70 0.83
10 9 $0.45\pi$ 0.71 0.65 0.80 0.71 0.65 0.80
@@
Table 5: Paired POD modes using the analytic signal of the $k^{th}$ POD mode, and the corresponding phase angle, correlation coefficient and $R^{2}_{k^{a},\widetilde{(k,j)}}$ of the phase averages. Superscripts u and v refer to the streamwise and transverse components, respectively.
@@
For each of the first ten POD modes, the paired POD mode corresponds to the analytic signal at a phase angle of $\mp\pi/2$. Consequently, the $j^{th}$ POD mode best matches the Hilbert transform of the corresponding $k^{th}$ POD mode. A pair of POD modes representing a structure propagating in the streamwise direction therefore requires that the POD modes be a mutual best match with phase angle shifts of $\pm 0.5\pi$. Based on this criterion, the POD mode pairs identified using the Hilbert transform are the first and second POD modes, the fourth and fifth POD modes, the sixth and eighth POD modes, and the ninth and tenth POD modes. These are the same pairings as those identified using the HPOD modes.
@@
Figure 11: Phase-averaged (a) streamwise and (c) transverse components of the first and second POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the first POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 5 for an animated version of this figure.
The streamwise component of the analytic signal of the first POD mode, shown in figure 11b, closely matches the phase average of the first two POD modes, shown in figure 11a. Both exhibit extended regions of alternating sign propagating downstream from the sphere. The primary difference occurs for $x/D<2$, where the structures in the POD modes are seen coming around the sphere, as illustrated in figure 6a(iv). Conversely, the assumed periodicity of the Hilbert transform couples these structures with those further downstream in the HPOD mode, as shown in figure 11b(iv). This discrepancy arises from applying the Hilbert transform to the first POD mode, as these structures exist in the second POD mode but not in the first. Consequently, the $R^{2}_{k^{a},\widetilde{(k,j)}}$ values differ: 0.89 for the analytic signal of the first mode when compared to the second mode, and 0.76 for the reverse comparison. This variation arises because non-propagating components in one mode are absent from the other but are propagated by the Hilbert transform. The variation between the transverse components of the phase-averaged first and second POD modes, shown in figure 11c, and the analytic signal of the first POD mode, shown in figure 11d, is considerably smaller than that observed in the streamwise components. Both exhibit symmetric regions of alternating sign that elongate and weaken in intensity as they propagate downstream. This behaviour is consistent with the transverse fluctuations generated by the flow around the sphere decaying as they travel downstream.
@@
Figure 12: Phase-averaged (a) streamwise and (c) transverse components of the fourth and fifth POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the fourth POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 6 for an animated version of this figure.
@@
The streamwise component of the analytic signal of the fourth POD mode, shown in figure 12b, retains more of the asymmetry of the phase-averaged fourth and fifth POD modes, shown in figure 12a, than is observed in the second HPOD mode, shown in figure 7b. The transverse component of the analytic signal of the fourth POD mode, shown in figure 12d, also retains the asymmetry present in the phase average of the fourth and fifth POD modes, shown in figure 12c. Although the bias of the Hilbert transform towards propagating structures is less pronounced in the analytic signal of the fourth POD mode than in the second HPOD mode, shown in figures 7b and 7d, it remains evident, particularly in figures 12b(iii) and 12d(iv).
@@
Figure 13: Phase-averaged (a) streamwise and (c) transverse components of the sixth and eighth POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the sixth POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 7 for an animated version of this figure.
@@
While the analytic signal of the sixth POD mode exhibits more symmetry than the phase average of the sixth and eighth POD modes, its influence on the turbulent structures is less pronounced than that observed in the third HPOD. The analytic signal of the sixth POD mode introduces the negative region at $x/D=2.5$ in the streamwise component shown in figure 13a(v), similar to that observed in the third HPOD mode shown in figure 8b(v), which is not present in the phase average of the sixth and eighth POD modes, shown in figure 13b(v). The analytic signal of the sixth POD mode also retains more of the transverse shifting in the structures in the transverse velocity, as seen in figure 13d, than the third HPOD mode, shown in figure 8d. Additionally, the intensity of the structures further downstream is preserved in the analytic signal of the sixth POD mode, as illustrated in figure 13b(i), where the same region exhibits significantly reduced intensity in the third HPOD mode, shown in figure 8b(i).
@@
Figure 14: Phase-averaged (a) streamwise and (c) transverse components of the ninth and tenth POD modes and the phase-averaged (b) streamwise and (d) transverse components of the phase average of the analytic signal of the ninth POD mode at $\phi=$ (i) 0, (ii) $\pi/6$, (iii) $\pi/3$, (iv) $\pi/2$, (v) $2\pi/3$, and (vi) $5\pi/6$. See supplementary movie 8 for an animated version of this figure.
@@
While there is a significant difference between the streamwise components of the analytic signal of the ninth POD mode, shown in figure 14b, and that of the phase average of the ninth and tenth POD modes, shown in figure 14a, for $x/D<2$, they become more similar further downstream of the sphere. The structures in the phase average of the ninth and tenth modes are smaller in scale for $x/D<2$ than they are for $x/D>2$, which is consistent with the breakdown of smaller-scale fluctuations in the near wake as they travel downstream while the larger-scale structures persist. Consequently, these structures are less significant to the Hilbert transform and therefore attenuated in the analytic signal. The differences between the transverse components of the analytic signal of the ninth POD mode, shown in figure 14d, and the phase average of the ninth and tenth POD modes, shown in figure 14c, are less pronounced. However, some variation remains in the upstream structures at $x/D>1.5$, as shown in figure 14c(ii) and 14d(ii), and figures 14c(iii) and 14d(iii), particularly at the end of the domain, which results from spectral leakage in the analytic signal caused by the non-periodicity of the POD modes.